# Automated Multi-Task Learning for Joint Disease Prediction on Electronic Health Records

### v1

@@ 3.2. Overview
(3) $\mathbf{g}_{(C,A)}=F(C,A),C\in\mathcal{C},A\in\mathcal{A},$
@@ 5                   Randomly sample $Q_{1}$ architectures from $\mathcal{A}$, denote the set as $\mathcal{A}^{C_{j}}$;
8                   Calculate the mean over top $Q_{2}$ gains from $\mathcal{G}$, denoted as ${g}^{C_{j}}$;
9
@@ 8                   Calculate the mean over top $Q_{2}$ gains from $\mathcal{G}$, denoted as ${g}^{C_{j}}$;
11            Build the sampling distribution over $\mathcal{C}^{T_{n}}$ as: $\mathbf{\Gamma}(\mathcal{C}^{T_{n}})=\{{\rm exp}({g}^{C_{j}}),\forall C_{j}\in\mathcal{C}^{T_{n}}\}$;
12             Sample a task combination $C^{*}$ from $\mathcal{C}^{T_{n}}$ based on $\mathbf{\Gamma}(\mathcal{C}^{T_{n}})$, and randomly sample an architecture $A^{*}$ from $\hat{\mathcal{A}}^{C^{*}}$;
@@ 17      Update $\mathbf{\Theta}$ by training the surrogate model on $\mathcal{D}$;
Algorithm 1 Active Sampling
3.3. Surrogate Model
3.3.1. Architecture Encoding
For encoding a given architecture $A$, we apply a graph encoder (Zhang et al., 2019) that is specifically designed for modeling DAGs, which is suitable for encoding the architectures in our search space. It can sequentially update the hidden states for the $P$ computation nodes in preceding order by aggregating information from all predecessors. For node $p$, we have:
(4) $\mathbf{h}_{p}={\rm Aggregate}(\mathbf{W}_{0}\cdot\mathbf{h}_{0},\mathbf{W}_{1}\cdot\mathbf{h}_{1},\cdots,\mathbf{W}_{p-1}\cdot\mathbf{h}_{p-1}),$
3.3.2. Task Combination Encoding
For encoding a given task combination $C$, we use the self attention mechanism (Vaswani et al., 2017) to model the high order interactions among the selected tasks in $C$. Specifically, we randomly initialize the embedding for all $N$ tasks, and for task combination $C$, we have:
(5) $\mathbf{z}={\rm Pool}({\rm SelfAttention}(\mathbf{u}_{1},\mathbf{u}_{2},\cdots,\mathbf{u}_{|C|})),$
where $\mathbf{u}\in\mathbb{R}^{d_{s}}$’s are corresponding embeddings for the selected tasks and $\mathbf{z}\in\mathbb{R}^{d_{s}}$ is final representation for task combination $C$. Additionally, we also use average pooling on top of the self attention layers to obatin $\mathbf{z}$.
Eventually, we apply a two layer MLP to fuse both architecture encoding $\mathbf{h}_{P}$ and task combination encoding $\mathbf{z}$, and output the predicted gains for all selected tasks $\hat{\mathbf{g}}_{(C,A)}\in\mathbb{R}^{|C|}$. We use the mean absolute error to supervise the surrogate model as follows:
(6) $\mathcal{L}(\hat{\mathbf{g}}_{(C,A)},\mathbf{g}_{(C,A)})=||\hat{\mathbf{g}}_{(C,A)}-\mathbf{g}_{(C,A)}||_{1},$
3.4. Active Sampling
In order to efficiently train the surrogate model defined in previous section, we develop an active sampling method to collect training samples. That is, our method progressively selects the samples that bring higher predicted gains for each task. In this way, the trained surrogate model is able to better facilitate the final selection of optimal task groupings and architectures, which also also improves the sample efficiency for achieving the expected performance gain.
3.4.1. Warm start
Firstly, we warmup the surrogate model by selecting a small number of samples from the search space. Specifically, we use the task combination that contains all $N$ tasks $C_{0}=\{T_{1},\cdots,T_{n}\}$ and randomly sample $Q_{0}$ architectures from $\mathcal{A}$. Then we conduct $Q_{0}$ MTL procedures to evaluate their transferring gains by training on $C_{0}$. In this way, we collect $Q_{0}$ training samples as the initial training set denoted as $\mathcal{D}$. And, we train the surrogate model on $D$, and denote the model parameters as $\mathbf{\Theta}$.
3.4.2. Active selection
@@
Then, we conduct $K_{1}$ rounds of active sampling. For each round, we iterate through all $N$ tasks. With respect to each task $T_{n}$, we build a distribution $\mathbf{\Gamma}$ over the the set of task combinations that contains $T_{n}$ based on the predicted gains for $T_{n}$. From the distribution, we sample one task combination $C^{*}$ that has a large probability to achieve higher gain for $T_{n}$. Then, we also select one architecture $A^{*}$ with high predicted gain for $T_{n}$ when combined with $C^{*}$. The selection of $C^{*}$ and $A^{*}$ is interdependent, and the details are introduced in Algorithm 1. In this way, we collect one sample $(C^{*},A^{*})$ to update the training set $\mathcal{D}$ with respect to each $T_{n}$. At the end of each round, we also update the surrogate model parameters $\mathbf{\Theta}$ with the updated $\mathcal{D}$. After $K_{1}$ rounds, we are able to obtain a well trained surrogate model for estimating the whole search space.
Input: Trained surrogate model $F(\cdot)$, $B$, $V$.
Output: Searched population $\mathcal{P}$.
2 for $v=1,2,\cdots,V$ do
3       Randomly select one pair $(C,A)$ from $\mathcal{P}$;
5       Estimate $\mathcal{P}^{\prime}$ and $\mathcal{P}$ using $F(\cdot)$;
7
8 end for
Algorithm 2 Greedy Search
3.5. Derivation
@@
We derive the final results using only the trained surrogate model. Because of the huge search space, it is still not practical to use brute force search to get the global optimum. Therefore, we propose to apply a greedy method to search for near-optimal solutions.
Given the budget $B$, we aim at searching for $B$ samples from the search space $\mathcal{C}\times\mathcal{A}$ such that the overall gain $G$ is maximized. We use a greedy algorithm to search for the optimal results as shown in Algorithm 2. Specifically, we first randomly initialize the population $\mathcal{P}$ that contains $B$ pairs of task combinations and architectures. Then, at every iteration, we randomly mutate one pair $(C,A)$ from the population and see whether the overall multi-task gain will increase. If so, we update $\mathcal{P}$ accordingly. After $K_{2}$ iterations, we can obtain a near-optimal solution. In practice, we also apply multiple initial populations to avoid getting stuck on local optima. Although we only get approximate solution, our method can already achieve significant improvement over baselines as shown in Section 4.2.
4. Experiments
4.1. Set Up
4.1.1. Dataset & Tasks
@@
We adopt MIMIC - IV dataset (Johnson et al., 2020) for our experiments, which is a publicly available database sourced from the electronic health record of the Beth Israel Deaconess Medical Center. Specifically, we extract the clinical time series data for the 56,908 ICU stays from the database as our input EHR data, with an average sequence length of 72.9. With respect to each ICU stay, we also extract 25 prediction tasks, including chronic, mixed, and acute care conditions. Each condition is associated with a binary label indicating whether the patient has the corresponding condition during the ICU stay.
To prepare our dataset, we adopt the data pre-processing pipeline outlined in Harutyunyan, et al. (Harutyunyan et al., 2019). Given that the original implementation [^3] https://github.com/YerevaNN/mimic3-benchmarks is designed for MIMIC-III (Johnson et al., 2016), we make specific modifications to tailor it for MIMIC-IV. The 25 labels are defined using the Clinical Classifications Software (CCS) for ICD-9 code [^4] https://www.cdc.gov/nchs/icd/icd9.htm. Consequently, we first map the ICD-10 codes [^5] https://www.cms.gov/medicare/coding-billing/icd-10-codes/2018-icd-10-cm-gem in the MIMIC-IV database to ICD-9 codes before generating the labels. After processing, we have the feature dimension $d_{e}$ as 76. We partition the dataset as train, validation and test sets with a ratio of $0.7:0.15:0.15$.
4.1.2. Baselines
@@
To compare the proposed method with existing work, we choose several state-of-art-baselines, including both hand-crafted and automated methods. Specifically, as described below, we include several human-designed EHR encoders to compare with the searched architecture we defined in Eq. (2). Also, we include one NAS method and one multi-task grouping method as the automated baselines. More importantly, we combine the multi-task grouping method with the NAS method and hand-crafted encoders to show the superiority of our joint optimization method.
@@
  • EHR encoders: We choose four models that are widely utilized for analyzing EHR time series, including LSTM (Hochreiter and Schmidhuber, 1997), Transformer (Vaswani et al., 2017), Retain (Choi et al., 2016b) and Adacare (Ma et al., 2020).
@@
  • NAS: We choose DARTS (Liu et al., 2019) as the NAS baseline, which proposes a differentiable search method for efficient architecture search. We apply it to our search space $\mathcal{A}$ to find better EHR encoders. It has also been adopted for searching better MTL architectures in most state-of-the-art works (Sun et al., 2020; Guo et al., 2020; Gao et al., 2020; Zhang et al., 2022).
@@
  • Multi-task grouping: MTG-Net (Song et al., 2022) is the current state-of-the-art multi-task grouping algorithm, which uses a meta learning approach to learn the high-order relationships among different tasks. We refer to this method as MTG in latter sections.
4.1.3. Evaluation Metric
@@
We use two widely used metrics for binary classification to evaluate our method and baselines: ROC (Area Under the Receiver Operating Characteristic curve) and AVP (Averaged Precision). During surrogate model training, we use AVP as the metric to compute multi-task gains as in Eq. (1).
Table 1. Performance of the single task backbone.

### v2

@@ 3.2 Overview
$\mathbf{g}_{(C,A)}=F(C,A),C\in\mathcal{C},A\in\mathcal{A},$ (3)
@@
3.3 Surrogate Model
Architecture Encoding. For encoding a given architecture $A$, we apply a graph encoder [29] that is specifically designed for modeling DAGs, which is suitable for encoding the architectures in our search space. It can sequentially update the hidden states for the $P$ computation nodes in preceding order by aggregating information from all predecessors. For node $p$, we have:
$\mathbf{h}_{p}={\rm Aggregate}(\mathbf{W}_{0}\cdot\mathbf{h}_{0},\mathbf{W}_{1}\cdot\mathbf{h}_{1},\cdots,\mathbf{W}_{p-1}\cdot\mathbf{h}_{p-1}),$ (4)
Task Combination Encoding. For encoding a given task combination $C$, we use the self attention mechanism [30] to model the high order interactions among the selected tasks in $C$. Specifically, we randomly initialize the embedding for all $N$ tasks, and for task combination $C$, we have:
$\mathbf{z}={\rm Pool}({\rm SelfAttention}(\mathbf{u}_{1},\mathbf{u}_{2},\cdots,\mathbf{u}_{|C|})),$ (5)
where $\mathbf{u}\in\mathbb{R}^{d_{s}}$’s are corresponding embeddings for the selected tasks and $\mathbf{z}\in\mathbb{R}^{d_{s}}$ is the final representation for task combination $C$. Additionally, we also use average pooling on top of the self attention layers to obatin $\mathbf{z}$.
Prediction. Eventually, we apply a two layer MLP to fuse both architecture encoding $\mathbf{h}_{P}$ and task combination encoding $\mathbf{z}$, and output the predicted gains for all selected tasks $\hat{\mathbf{g}}_{(C,A)}\in\mathbb{R}^{|C|}$. We use the mean absolute error to supervise the surrogate model as follows:
$\mathcal{L}(\hat{\mathbf{g}}_{(C,A)},\mathbf{g}_{(C,A)})=||\hat{\mathbf{g}}_{(C,A)}-\mathbf{g}_{(C,A)}||_{1},$ (6)
3.4 Progressive Sampling
In order to efficiently train the surrogate model defined in previous section, we develop a progressive sampling method to collect training samples. Start with an empty training set and a random initialized surrogate model, we progressively sample more points from the search space $\mathcal{C}\times\mathcal{A}$, and then use them to train the surrogate model. Specifically, we include two stages for the surrogate model training:
Warm start. Firstly, we warmup the surrogate model by selecting a small number of samples from the search space. Specifically, we use the task combination that contains all $N$ tasks $C_{0}=\{T_{1},\cdots,T_{n}\}$ and randomly sample $Q_{0}$ architectures from $\mathcal{A}$. Then we conduct $Q_{0}$ MTL procedures to evaluate their gains by training on $C_{0}$. In this way, we collect $Q_{0}$ training samples as the initial training set denoted as $\mathcal{D}$. We further train the surrogate model on $D$, and denote the model parameters as $\mathbf{\Theta}$.
1
@@ 7                   Randomly sample $Q_{1}$ architectures from $\mathcal{A}$, denote the set as $\mathcal{A}^{C_{j}}$;
10                   Calculate the mean over top $Q_{2}$ gains from $\mathcal{G}$, denoted as ${\mu}^{C_{j}}$;
11                   Calculate the variance over top $Q_{2}$ gains from $\mathcal{G}$, denoted as ${\sigma}^{C_{j}}$;
12
@@ 11                   Calculate the variance over top $Q_{2}$ gains from $\mathcal{G}$, denoted as ${\sigma}^{C_{j}}$;
14             Compute the acquisition values over $\mathcal{C}^{T_{n}}$ as: $\mathbf{\Gamma}(\mathcal{C}^{T_{n}})=\{{\mu}^{C_{j}}+\lambda\cdot{\sigma}^{C_{j}},\forall C_{j}\in\mathcal{C}^{T_{n}}\}$;
15             Sample a task combination $C^{*}$ from $\mathcal{C}^{T_{n}}$ that has highest value in $\mathbf{\Gamma}(\mathcal{C}^{T_{n}})$, and randomly sample an architecture $A^{*}$ from $\hat{\mathcal{A}}^{C^{*}}$;
@@ 20      Update $\mathbf{\Theta}$ by training the surrogate model on $\mathcal{D}$;
23
Algorithm 1 Progressive Selection
@@
Progressive selection. Then, we progressively select more points and train the surrogate model as introduced in Algorithm 1. Totally, we conduct $K_{1}$ rounds of sampling. For each round, we iterate through all $N$ tasks. With respect to one task $T_{n}$, we build the acquisition function $\mathbf{\Gamma}$ over the set of task combinations that contains $T_{n}$ based on the predicted gains for $T_{n}$. Then, we select one task combination $C^{*}$ that have highest value. We apply Upper Confidence Bound [31] as the acquisition function that considers both exploration and exploitation by explicitly estimating the mean and variance of predicted gains (line 11 marked by blue). Besides that, we would also like to see the effect of exploration vs exploitation. so we try out different settings of $\mathbf{\Gamma}$. Specifically, we propose three variants of AutoDP, namely AutoDPμ+σ, AutoDPμ and AutoDPσ, which corresponds to the original setting, including only $\mu$ or only $\sigma$ in $\mathbf{\Gamma}$. In this way, we can compare the results with pure exploration and pure exploitation during sampling, and find out the optimal strategy for AutoDP. Moreover, we also select one architecture $A^{*}$ with high predicted gain for $T_{n}$ when combined with $C^{*}$. The selection of $C^{*}$ and $A^{*}$ is interdependent, and the details are introduced in Algorithm 1. In this way, we collect one sample $(C^{*},A^{*})$ to update the training set $\mathcal{D}$ with respect to each $T_{n}$. At the end of each round, we also update the surrogate model parameters $\mathbf{\Theta}$ with the updated $\mathcal{D}$. After $K_{1}$ rounds, we are able to obtain a well trained surrogate model for estimating the whole search space.
3.5 Derivation
@@
We derive the final results using the trained surrogate model. Due to the huge search space, it is still not practical to use brute force search to get the global optimum. Therefore, we propose to apply a greedy method to search for near-optimal solutions. We introduce the detailed procedure in Appendix C. The high level idea is that we first randomly initialize the configuration, and then gradually improve its multi-task gain by random mutation and greedy selection. After a reasonably large number of iterations, we are able to attain the approximate solution that can perform well as shown in Section 4.2.
4 Experiments
4.1 Set Up
@@
Dataset & Tasks. We adopt MIMIC - IV dataset [25] for our experiments, which is a publicly available database sourced from the electronic health record of the Beth Israel Deaconess Medical Center. Specifically, we extract the clinical time series data for the 56,908 ICU stays from the database as our input EHR data, with an average sequence length of 72.9. With respect to each ICU stay, we also extract 25 prediction tasks (listed in Table 3), including chronic, mixed, and acute care conditions. Each condition is associated with a binary label indicating whether the patient has the corresponding condition during the ICU stay.
@@
Baselines. To compare the proposed method with existing work, we choose several state-of-art-baselines, including both hand-crafted and automated methods. Specifically, as described below, we include several human-designed EHR encoders to compare with the searched architecture we defined in Eq. (2). Also, we include one NAS method and one multi-task grouping method as the automated baselines. More importantly, we combine the multi-task grouping method with the NAS method and hand-crafted encoders to show the superiority of our joint optimization method.
@@
  • EHR encoders: We choose four models that are widely utilized for analyzing EHR time series, including LSTM [32], Transformer [30], Retain [22] and Adacare [1].
@@
  • NAS: We choose DARTS [23] as the NAS baseline, which is a differentiable search method for efficient architecture search. We apply it to our search space $\mathcal{A}$ to find better EHR encoders. Several state-of-the-art works in other domains have also used it to find MTL architectures [16, 17, 18, 21].
@@
  • Multi-task grouping: MTG-Net [13] is the current state-of-the-art multi-task grouping algorithm, which uses a meta learning approach to learn the high-order relationships among different tasks. We refer to this method as MTG in latter sections.
@@
Evaluation Metric. We use two widely used metrics for binary classification to evaluate our method and baselines: ROC (Area Under the Receiver Operating Characteristic curve) and AVP (Averaged Precision). During surrogate model training, we use AVP as the metric to compute multi-task gains as in Eq. (1), since it is a more suitable choice for considering the class imbalance.
Please also refer to Appendix A for the implementation details.
4.2 Performance Evaluation
We show our results in Table 1. Each experiment is run five times and the average of the runs are reported. We run three settings: Task @ 5, Task @ 10 and Task @ 25, which refers to using the first 5 tasks, 10 tasks and 25 tasks respectively. Since grouping all 25 tasks takes a long time to run, we include two small settings that only have the first 5 or 10 tasks in Table 3 for grouping. Our results demonstrate our hypotheses: (a) applying Retain, Adacare, and DARTS improves over the single-task setting, (b) applying different NAS models for each group further improves the performance, and finally, (c) AutoDP provides the best results in terms of averaged per-task gain for ROC and AVP, a significant improvement over existing MTL frameworks for EHR data.
Table 1: Performance comparison in terms of averaged per-task gain over single task backbone (All results are in the form of percentage values %).
Settings Included Tasks Tasks @ 5 Tasks @ 10 Tasks @ 25
Metric ROC AVP ROC AVP ROC AVP
One model for all tasks LSTM +0.09 +0.18 +1.06 +3.22 +1.83 +7.46
Transformer +0.97 +4.82 +1.41 +4.14 +1.75 +7.45
Retain +0.46 +1.80 +0.66 +0.75 +1.41 +5.88
Adacare +1.03 +5.21 +1.32 +4.05 +1.68 +6.94
DARTS +1.28 +5.01 +2.01 +6.87 +1.87 +7.71
Task Grouping + One model for each group MTG+LSTM +0.51 +2.10 +0.65 +1.87 +1.74 +7.40
MTG+Transformer +0.91 +3.64 +1.20 +3.95 +1.79 +9.15
MTG+Retain +0.55 +3.11 +1.51 +5.20 +1.54 +8.87
MTG+Adacare +1.25 +5.78 +1.44 +4.63 +1.75 +7.84
MTG+DARTS +1.47 +6.41 +2.02 +6.65 +2.41 +11.76
Variants of AutoDP AutoDPμ +1.49 +7.12 +2.08 +7.53 +2.68 +12.70
AutoDPσ +1.95 +7.68 +2.49 +8.45 +2.62 +13.37
AutoDPμ+σ +1.69 +7.74 +2.55 +8.81 +2.80 +13.43
(std) $\pm$ 0.08 $\pm$ 0.25 $\pm$ 0.13 $\pm$ 0.29 $\pm$ 0.12 $\pm$ 0.33
(p-value) 0.045 0.029 0.036 0.045 0.027 0.032
First, without considering task grouping, we train one shared model to predict for all tasks in three settings and compute the multi-task gains for them. Results show that this setting only provides minimal improvement over single task training. Note that the automated method (DARTS) performs better than other hand-crafted methods. We also see that sequential optimization over task grouping and architecture search (MTG+DARTS) performs better than MTG + other hand-crafted encoders.
Finally, we see that the three variants of AutoDP performed better than the other methods. Among them, AutoDPμ+σ performs the best, which means the balance of exploration and exploitation is the most effective strategy for training the surrogate model. For the last method, we also report the standard deviations and p-values of statistical tests (compared to MTG+DARTS), which justifies that the improvement is significant. The runtime is approximately as the same for MTG+DARTS and AutoDP and thus this is a fair comparison.
We also include more results for individual task gains of AutoDP, please refer to Appendix B.
4.3 Hyperparameter & Complexity Analysis
We analyze the effect of two vital hyperparameters of our method: $K_{1}$ and $B$, since they are the crucial parameters that largely define the complexity of our method during searching and inference respectively. We choose the setting of Task @ 25 for a comprehensive analysis of all tasks. We try out different values and report the corresponding performance gain (AVP) in Figure 2.
Figure 2: Analysis for the number of progressive sampling rounds $K_{1}$ and the budget of task groups $B$ under the setting of Task @ 25.
First, $K_{1}$ determines the number of training samples collected during searching. Given that each sample invokes an MTL procedure, it constitutes the major portion of the search cost. Therefore, our goal is to find an optimal value for $K_{1}$, striking a balance between cost-effectiveness and achieving commendable performance. We notice that the performance change plateaus after $K_{1}$ reaches 25. That is, the surrogate model effectively learns the distribution of the search space after consuming $25\times 25$ training samples during active selection (25 samples per round). Consequently, we can empirically decide to halt the iteration at this point.
Second, $B$ determines the number of task groups for the final configuration, which indicates the number of MTL models required for achieving the expected performance gain after searching. We also observe similar phenomenon that the performance becomes stable after $B$ reaches 12. We could also choose the optimal value for $B$ accordingly.
4.4 Ablation Study
Table 2: Ablation results in terms of AVP.
Settings Task @ 5 Task @ 10 Task @ 25
AutoDP +7.74 +8.81 +13.43
Random sampling +6.75 +7.04 +11.30
Random search +6.89 +7.15 +12.04
Disease grouping +6.29 +6.99 +8.61
We further analyze the effect of several components within AutoDP, including progressive sampling (Section 3.4), greedy search (Section 3.5), and task grouping as a whole. We replace these components with naive or human intuition-inspired baselines and report the performances in Table 2. Removing any of the components from the original framework leads to noticeable performance decreases, demonstrating the effectiveness of the designed components.
Specifically, we replace progressive sampling and greedy search with purely random methods, referred to as Random Sampling and Random Search. In all three settings, performance generally decreases, highlighting the contributions of these components of AutoDP.
Additionally, we use disease-based grouping (Appendix D) to first assign tasks into different groups based on their medical relevance and then employ DARTS to search for the model architecture for each group. This allows us to analyze the effectiveness of automated task grouping. By comparing disease-based grouping with the searched configurations (Appendix E), we observe that AutoDP does not strictly follow medical classifications for task grouping but achieves significant performance improvements over disease-based grouping. This indicates the necessity of using an automated search algorithm to find the optimal task grouping, which surpasses human intuition.
5 Conclusions
In this paper, we propose AutoDP, an automated multi-task learning framework for joint disease prediction on EHR data. Compared to existing work, our method largely improves the design of task grouping and model architectures by reducing human interventions. Experimental results on real-world EHR data demonstrate that the proposed framework achieves significant improvement over existing state-of-the-art methods, while maintaining a feasible search cost.
Appendix
We include the following sections in Appendix, providing the additional details of the proposed framework, additional experimental results, and a discussion about limitation and future work.
Appendix A Implementation
To prepare our dataset, we adopt the data pre-processing pipeline outlined in Harutyunyan, et al. [27]. Given that the original implementation [^2] https://github.com/YerevaNN/mimic3-benchmarks is designed for MIMIC-III [33], we make specific modifications to tailor it for MIMIC-IV. The 25 labels are defined using the Clinical Classifications Software (CCS) for ICD-9 code [^3] https://www.cdc.gov/nchs/icd/icd9.htm. Consequently, we first map the ICD-10 codes [^4] https://www.cms.gov/medicare/coding-billing/icd-10-codes/2018-icd-10-cm-gem in the MIMIC-IV database to ICD-9 codes before generating the labels. After processing, we have the feature dimension $d_{e}$ as 76. We partition the dataset as train, validation and test sets with a ratio of $0.7:0.15:0.15$.
We implement the framework using the PyTorch framework and run it on an NVIDIA A100 GPU. Given the dataset we have, we first train a vanilla LSTM for every task independently, and report the backbone performance in Table 3, which can be further used to compute multi-task gains. For the proposed method, we run three settings of experiments: Task @ 5, Task @ 10 and Task @ 25, which refers to using the first 5 tasks, 10 tasks and 25 tasks respectively. For different settings, we use specific hyperparameters as shown in Table 4. Besides that, we define the candidate operation set $\mathcal{O}$ as {Identity, Zero, FFN, RNN , Attention}, which includes widely used operations for processing EHR time series. Among them, Identity means maintaining the output identical to the input. Zero means setting all the values of the input feature to 0. Attention and FFN represents one self-attention layer and one feed-forward layer respectively, which are the same as in Transformer [30]. RNN is one recurrent layer, and we adopt LSTM [32] in our framework. For all the MTL procedures and baseline training, we apply the batch size of 64 and learning rate of $43e-$ 4. For training the surrogate model, we use the batch size of 5 and learning rate of $55e-$ 5. During searching, we compute all multi-task gains on the validation set for guiding the surrogate model training. After we obtain the optimal configuration, we train the searched models and report their multi-task gains on the test set.
Table 3: Performance of the single task backbone.